# Interpolation theorems for conjugations and applications

### v1

@@ 2.1. Main results
Let $N\in{\mathcal{B}}(H)$ be normal and $M$ a subspace of $H$. Then, $M$ is hyperinvariant for $N$ if and only if $CM\subseteq M$ for every conjugation $C$ on $H$ satisfying $CNC=N^{*}$.
The proof of the necessity is similar to the one in [20]. We include it here for the sake of completeness.
Proof of Proposition 2.5.
Assume that $M$ is hyperinvariant for $N$ and denote by $P\in{\mathcal{B}}(H)$ the orthogonal projection onto $M$. Let $C$ be a conjugation on $N$ satisfying $CNC=N^{*}$; equivalently, $CN=N^{*}C$. Clearly, we need to show that $CP=PC$. According to [25, Proposition 6.9], there exists a Borel subset $\Delta\subseteq\mathbb{C}$ such that $P=E_{N}(\Delta)$. Consequently, Lemma 2.11 implies that $CP=CE_{N}(\Delta)=E_{N}(\Delta)C=PC$.
For the sufficiency, we need to show that for each $T\in\{N\}^{\prime}$, the commutant of $N$, $TM\subseteq M$. Additionally, since $\{N\}^{\prime}$ is a von Neumann algebra, it is spanned by its unitary operators ([9, Page 61, Proposition 13.3 (b)]). Consequently, we can assume without loss of generality that $T$ is unitary. Let $x\in M$ be non-zero, and put $y=Tx$. Then, according to Theorem 2.4, there exists a conjugation $C$ on $H$ satisfying $CNC=N^{*}$ and $Cx=y$. Hence, by hypothesis, $Tx=y=Cx\in M$. Since $x$ was arbitrary, we conclude that $TM\subseteq M$. ∎
The following example shows that the result in Proposition 2.5 cannot be extended to an arbitrary operator in ${\mathcal{B}}(H)$.
@@
Let $x$ and $y$ be orthonormal vectors in $H$, and put $T=A+{\rm i}B$ where $A=x\otimes x$ and $B=(x+y)\otimes(x+y)$. Clearly, $A$ and $B$ are self-adjoint operators; moreover, $A=\frac{1}{2}(T+T^{*})$. It follows that if $C$ is a conjugation on $H$ such that $CTC=T^{*}$, then $CAC=A$. Hence, if $M={\rm span}\{x\}$, then $CM=M$. Therefore, the subspace $M$ satisfies the necessary condition in Proposition 2.5; however, $T\in\{T\}^{\prime}$ and $TM\nsubset M$.
Given a conjugation $C$ on $H$, an operator $T\in{\mathcal{B}}(H)$ is said to be $C$-normal if $CTT^{*}C=T^{*}T$. When $T$ is $C$-normal for some conjugation $C$, we say that $T$ is conjugate-normal. It is easy to verify that if $T$ is $C$-symmetric or $C$-skew symmetric, then it is also $C$-normal. Consequently, conjugate-normality is considered a more general type of symmetry. For further details and properties of such operators, the interested reader is referred to [27].
@@
Let $T\in{\mathcal{B}}(H)$. The following statements are equivalent:
@@
    For every non-zero $x\in H$, there exists a conjugation $C$ on $H$ such that $T$ is $C$-symmetric and $Cx=x$;
@@
    For every non-zero $x\in H$, there exists a conjugation $C$ on $H$ such that $T$ is $C$-normal and $Cx=x$.
@@
(iii)$\implies$(ii). Let $x\in H$ be non-zero. Then, by hypothesis, there exists a conjugation $C$ on $H$ such that $Cx=x$ and $CTT^{*}C=T^{*}T$. It follows that
@@
2.2. Proofs of Theorems 2.1, 2.3 and 2.4
The proofs require some preparation and will be provided at the end of this section. In what follows, we denote the set of all complex polynomials in two variables by $\mathbb{C}[x,y]$. We begin with the following technical lemma.
@@
Let $N\in{\mathcal{B}}(H)$ and $M\in\mathcal{B}(K)$ be normal operators, and let $x,y\in H$ and $h,k\in K$ be vectors such that
$\langle E_{N}(\Delta)x,y\rangle=\langle E_{M}(\Delta)h,k\rangle\quad\mbox{for every Borel subset $\Delta\subseteq\mathbb{C}$}.$
Then, for every $p\in\mathbb{C}[x,y]$, we have $\langle p(N,N^{*})x,y\rangle=\langle p(M,M^{*})h,k\rangle.$
@@
Let $p\in\mathbb{C}[x,y]$, and put $f(z)=p(z,\overline{z})\chi_{K}(z)$ where $K=\sigma(N)\cup\sigma(M)$ and $\chi_{K}$ is the characteristic function of $K$. Then, $f$ is a bounded Borel function on $\mathbb{C}$; furthermore, $f(N)=p(N,N^{*})$ and $f(M)=p(M,M^{*})$. This implies that
$\langle p(N,N^{*})x,y\rangle=\langle f(N)x,y\rangle=\int f{\,\rm d}\langle E_{N}(.)x,y\rangle$
@@
$\langle p(M,M^{*})h,k\rangle=\langle f(M)h,k\rangle=\int f{\,\rm d}\langle E_{M}(.)h,k\rangle.$
As the complex-valued measures $\langle E_{N}(.)x,y\rangle$ and $\langle E_{M}(.)h,k\rangle$ are presumed equal by hypothesis, we deduce $\int f{\,\rm d}\langle E_{N}(.)x,y\rangle=\int f{\,\rm d}\langle E_{M}(.)h,k\rangle$. Consequently, $\langle p(N,N^{*})x,y\rangle=\langle p(M,M^{*})x,y\rangle.$
Given a normal operator $N\in{\mathcal{B}}(H)$ and a family of vectors $X\subset H$, we denote
$L_{N,X}=\bigvee\big{\{}p(N,N^{*})h:p\in\mathbb{C}[x,y]\mbox{ and }h\in X\big{\}}$
where $\vee$ denotes the closed linear span. It is easy to see that $L_{N,X}$ is a reducing subspace of $N$. The restriction of $N$ to $L_{N,X}$ will be denoted by $N_{X}$.
@@
Let $N\in{\mathcal{B}}(H)$ and $M\in\mathcal{B}(K)$ be normal operators, and let $X=\{x_{i}\}_{i\in I}\subseteq H$ and $Y=\{y_{i}\}_{i\in I}\subseteq K$ be two families of vectors such that
$\langle E_{N}(\Delta)x_{i},x_{j}\rangle=\langle E_{M}(\Delta)y_{j},y_{i}\rangle\quad\mbox{for all Borel subsets $\Delta\subseteq\mathbb{C}$ and $i,j\in I$}.$
Then, there exists a conjugate-linear surjective isometry $W:L_{N,X}\to L_{M,Y}$ such that
$Wx_{i}=y_{i}\quad\mbox{and}\quad Wp(N_{X},N_{X}^{*})=p(M_{Y},M_{Y}^{*})^{*}W\quad\mbox{for all $i\in I$ and $p\in\mathbb{C}[x,y]$}.$
@@
According to Lemma 2.8, for every polynomial $p\in\mathbb{C}[x,y]$ and all $i,j\in I$, we have
(2.1) $\left\langle p(N,N^{*})x_{i},x_{j}\right\rangle=\left\langle p(M,M^{*})y_{j},y_{i}\right\rangle.$
@@
$H_{0}={\rm span}\big{\{}p(N,N^{*})x_{i}:p\in\mathbb{C}[x,y],i\in I\big{\}}$
@@
$K_{0}={\rm span}\big{\{}p(M,M^{*})y_{i}:p\in\mathbb{C}[x,y],i\in I\big{\}},$
and let $W:H_{0}\to K_{0}$ be the conjugate-linear transformation given by
$W(p(N,N^{*})x_{i})=p(M,M^{*})^{*}y_{i}\quad\mbox{for all $p\in\mathbb{C}[x,y]$and $i\in I$.}$
@@
We will first show that $W$ is well-defined. Arbitrarily choose $h\in H_{0}$; then, there exists a finite subset $J$ of $I$ and polynomials $p_{i}\in\mathbb{C}[x,y]$, for $i\in J$, such that $h=\sum_{i\in J}p_{i}(N,N^{*})x_{i}$. Then,
$\|h\|^{2}$ $=\left\langle\sum_{i\in J}p_{i}(N,N^{*})x_{i},\sum_{j\in J}p_{j}(N,N^{*})x_{j}\right\rangle$
$=\sum_{i\in J}\sum_{j\in J}\left\langle p_{i}(N,N^{*})x_{i},p_{j}(N,N^{*})x_{j}\right\rangle$
$=\sum_{i\in J}\sum_{j\in J}\left\langle p_{j}(N,N^{*})^{*}p_{i}(N,N^{*})x_{i},x_{j}\right\rangle$
[…]
$=\sum_{i\in J}\sum_{j\in J}\left\langle p_{j}(M,M^{*})^{*}y_{j},p_{i}(M,M^{*})^{*}y_{i}\right\rangle$
$=\left\langle\sum_{j\in J}p_{j}(M,M^{*})^{*}y_{j},\sum_{i\in J}p_{i}(M,M^{*})^{*}y_{i}\right\rangle=\|Wh\|^{2}.$
@@
Hence, if $h,k\in H_{0}$ are such that $\|h-k\|=0$, we get $\|Wh-Wk\|=\|W(h-k)\|=\|h-k\|=0$. Therefore, $W$ is a well-defined conjugate-linear isometry. Moreover, by its construction, $W$ is surjective. By density of $H_{0}$ and $K_{0}$ in $L_{N,X}$ and $L_{M,Y}$, respectively, it follows that $W$ can be extended to a conjugate-linear surjective isometry from $L_{N,X}$ to $L_{M,Y}$, which we also denote by $W$. Furthermore, it can be seen that $Wx_{i}=y_{i}$ for every $i\in I$.
Now, fix $p\in\mathbb{C}[x,y]$ and let us check that $Wp(N_{X},N_{X}^{*})=p(M_{Y},M_{Y}^{*})^{*}W$. Let $q\in\mathbb{C}[x,y]$, $i\in I$, and put $h=q(N,N^{*})x_{i}$. Then,
$Wp(N_{X},N_{X}^{*})h=Wp(N,N^{*})h$ $=Wp(N,N^{*})q(N,N^{*})x_{i}=p(M,M^{*})^{*}q(M,M^{*})^{*}y_{i}$
$=p(M,M^{*})^{*}Wq(N,N^{*})x_{i}=p(M_{Y},M_{Y}^{*})^{*}Wh.$

### v3

@@ 2.1. Main results
Let $N\in{\mathcal{B}}(\mathcal{H})$ be normal and $M$ a subspace of $\mathcal{H}$. The following statements are equivalents:
  1. (i)
    $M$ is hyperinvariant for $N$.
  2. (ii)
    $TM\subseteq M$ for every conjugate-linear operator $T$ on $\mathcal{H}$ satisfying $TN=N^{*}T$.
  3. (iii)
    $CM\subseteq M$ for every conjugation $C$ on $\mathcal{H}$ satisfying $CNC=N^{*}$.
Proof.
(i)$\implies$(ii). Assume that $M$ is hyperinvariant for $N$, and let $T$ be a conjugate-linear operator on $\mathcal{H}$ such that $TN=N^{*}T$. Clearly, it suffices to show that $TP_{M}=P_{M}T$. According to [26, Proposition 6.9], there exists a Borel subset $\Delta\subseteq\mathbb{C}$ such that $P_{M}=E_{N}(\Delta)$. Consequently, Lemma 2.15 implies that $TP_{M}=TE_{N}(\Delta)=E_{N}(\Delta)T=P_{M}T$.
(ii)$\implies$(iii) If $C$ is a conjugation on $\mathcal{H}$ such that $CNC=N^{*}$, and since $C^{-1}=C$, it follows that $CN=N^{*}C$, which leads to the desired containment.
(iii)$\implies$(i). We need to show that for each $T\in\{N\}^{\prime}$, the commutant of $N$, $TM\subseteq M$. Additionally, since $\{N\}^{\prime}$ is a von Neumann algebra, it is spanned by its unitary operators ([10, p. 61, Proposition 13.3 (b)]). Consequently, we can assume without loss of generality that $T$ is unitary. Let $x\in M$ be non-zero, and put $y=Tx$. Then, according to Theorem 2.8, there exists a conjugation $C$ on $\mathcal{H}$ satisfying $CNC=N^{*}$ and $Cx=y$. Hence, by hypothesis, $Tx=y=Cx\in M$. Since $x$ was arbitrary, we conclude that $TM\subseteq M$. ∎
The following example shows that the result in Proposition 2.10 cannot be extended to an arbitrary operator in ${\mathcal{B}}(\mathcal{H})$.
@@
Let $x$ and $y$ be orthonormal vectors in $\mathcal{H}$, and put $T=A+{\rm i}B$ where $A=x\otimes x$ and $B=(x+y)\otimes(x+y)$. Clearly, $A$ and $B$ are self-adjoint operators; moreover, $A=\frac{1}{2}(T+T^{*})$. It follows that if $C$ is a conjugation on $\mathcal{H}$ such that $CTC=T^{*}$, then $CAC=A$. Hence, if $M=\vee\{x\}$, then $CM=M$. Therefore, the subspace $M$ satisfies condition (iii) in Proposition 2.10; however, $T\in\{T\}^{\prime}$ and $TM\nsubset M$.
Given a conjugation $C$ on $\mathcal{H}$, an operator $T\in{\mathcal{B}}(\mathcal{H})$ is said to be $C$-normal if $CTT^{*}C=T^{*}T$. When $T$ is $C$-normal for some conjugation $C$, we say that $T$ is conjugate-normal. It is easy to verify that if $T$ is $C$-symmetric or $C$-skew symmetric, then it is also $C$-normal. Consequently, conjugate-normality is considered a more general type of symmetry. For further details and properties of such operators, the interested reader is referred to [28].
@@
Let $T\in{\mathcal{B}}(\mathcal{H})$. The following statements are equivalent:
@@
    For every unit vector $x\in\mathcal{H}$, there exists a conjugation $C$ on $\mathcal{H}$ such that $T$ is $C$-symmetric and $Cx=x$.
@@
    For every unit vector $x\in\mathcal{H}$, there exists a conjugation $C$ on $\mathcal{H}$ such that $T$ is $C$-normal and $Cx=x$.
@@
(iii)$\implies$(i). Let $x\in\mathcal{H}$ be a unit vector. Then, by hypothesis, there exists a conjugation $C$ on $\mathcal{H}$ such that $Cx=x$ and $CTT^{*}C=T^{*}T$. It follows that
@@
2.2. Proof of Theorems 2.1, 2.4, and 2.8
The proofs require some preparation and will be provided at the end of this section. In what follows, we denote the set of all complex polynomials in $n$ variables by $\mathcal{P}_{n}(\mathbb{C})$. We begin with the following technical lemma.
@@
Let $\{N_{i}\}_{1\leq i\leq n}\subset{\mathcal{B}}(\mathcal{H})$ and $\{M_{i}\}_{1\leq i\leq n}\subset\mathcal{B}(\mathcal{K})$ be two finite sets of mutually commuting normal operators, and let $x,y\in\mathcal{H}$ and $h,k\in\mathcal{K}$ be vectors such that
$\left\langle\prod_{i=1}^{n}\left(E_{N_{i}}(\Delta_{i})\right)x,y\right\rangle=\left\langle\prod_{i=1}^{n}\left(E_{M_{i}}(\Delta_{i})\right)h,k\right\rangle\quad\mbox{for every Borel subset $\Delta_{i}\subseteq\mathbb{C}$}.$
Then, for every $p\in\mathcal{P}_{2n}(\mathbb{C})$, we have
$\left\langle p(N_{1},N_{1}^{*},\dots,N_{n},N_{n}^{*})x,y\right\rangle=\left\langle p(M_{1},M_{1}^{*},\dots,M_{n},M_{n}^{*})h,k\right\rangle.$
@@
Let $p\in\mathcal{P}_{2n}(\mathbb{C})$. By the commutativity assumption, we may suppose that $p(\alpha_{1},\lambda_{1},\dots,\alpha_{n},\lambda_{n})=\prod_{j=1}^{n}\alpha_{i}^{r_{i}}\lambda_{i}^{s_{i}}$ where $r_{i},s_{i}\in\mathbb{N}$. For each $i\in\{1,\dots,n\}$, put $f_{i}(z)=z^{r_{i}}\overline{z}^{s_{i}}\chi_{K_{i}}(z)$ where $K_{i}=\sigma(N_{i})\cup\sigma(M_{i})$ and $\chi_{K_{i}}$ is the characteristic function of $K_{i}$. Then, for every $i\in\{1,\dots,n\}$, $f_{i}$ is a bounded Borel function on $\mathbb{C}$. Additionally,
(2.1) $f_{i}(N_{i})=N_{i}^{r_{i}}N_{i}^{*s_{i}}\quad\mbox{and}\quad f_{i}(M_{i})=M_{i}^{r_{i}}M_{i}^{*s_{i}}\quad\mbox{for every }i\in\{1,\dots,n\}.$
Fix Borel subsets $\{\Delta_{i}\}_{2\leq i\leq n}\subseteq\mathbb{C}$ and denote by $\mu_{1}$ and $\nu_{1}$ the complex-valued measures on Borel subsets of $\mathbb{C}$ given by
$\mu_{1}(\Delta)=\left\langle E_{N_{1}}(\Delta)\prod_{i=2}^{n}\left(E_{N_{i}}(\Delta_{i})\right)x,y\right\rangle\quad\mbox{and}\quad\nu_{1}(\Delta)=\left\langle E_{M_{1}}(\Delta)\prod_{i=2}^{n}\left(E_{M_{i}}(\Delta_{i})\right)h,k\right\rangle.$
Then,
$\left\langle N_{1}^{r_{1}}N_{1}^{*s_{1}}\prod_{i=2}^{n}\left(E_{N_{i}}(\Delta_{i})\right)x,y\right\rangle=\left\langle f_{1}(N_{1})\prod_{i=2}^{n}\left(E_{N_{i}}(\Delta_{i})\right)x,y\right\rangle=\int f_{1}{\,\rm d}\mu_{1}$
@@
$\left\langle M_{1}^{r_{1}}M_{1}^{*s_{1}}\prod_{i=2}^{n}\left(E_{M_{i}}(\Delta_{i})\right)h,k\right\rangle=\left\langle f_{1}(M_{1})\prod_{i=2}^{n}\left(E_{M_{i}}(\Delta_{i})\right)h,k\right\rangle=\int f_{1}{\,\rm d}\nu_{1}.$
As the measures $\mu_{1}$ and $\nu_{1}$ are presumed equal by hypothesis, we deduce
$\left\langle N_{1}^{r_{1}}N_{1}^{*s_{1}}\prod_{i=2}^{n}\left(E_{N_{i}}(\Delta_{i})\right)x,y\right\rangle=\left\langle M_{1}^{r_{1}}M_{1}^{*s_{1}}\prod_{i=2}^{n}\left(E_{M_{i}}(\Delta_{i})\right)h,k\right\rangle.$
By the Fuglede Theorem, we obtain $\mu_{2}(\Delta_{2})=\nu_{2}(\Delta_{2})$ where
$\mu_{2}(\Delta_{2}):=\left\langle E_{N_{2}}(\Delta_{2})N_{1}^{r_{1}}N_{1}^{*s_{1}}\prod_{i=3}^{n}\left(E_{N_{i}}(\Delta_{i})\right)x,y\right\rangle$
and
$\nu_{2}(\Delta_{2}):=\left\langle E_{M_{2}}(\Delta_{2})M_{1}^{r_{1}}M_{1}^{*s_{1}}\prod_{i=3}^{n}\left(E_{M_{i}}(\Delta_{i})\right)h,k\right\rangle.$
Since $\Delta_{2}$ was arbitrary, the complex-valued measures $\mu_{2}$ and $\nu_{2}$ are equal. Integrating $f_{2}$ with respect to these measures and taking into account (2.1), we obtain
$\left\langle N_{2}^{r_{2}}N_{2}^{*s_{2}}N_{1}^{r_{1}}N_{1}^{*s_{1}}\prod_{i=3}^{n}\left(E_{N_{i}}(\Delta_{i})\right)x,y\right\rangle=\left\langle M_{2}^{r_{2}}M_{2}^{*s_{2}}M_{1}^{r_{1}}M_{1}^{*s_{1}}\prod_{i=3}^{n}\left(E_{M_{i}}(\Delta_{i})\right)h,k\right\rangle.$
The desired equality will follow by continuing the previous process. ∎
@@
Let $\mathcal{A}=\{A_{k}\}_{k\in K}\subset{\mathcal{B}}(\mathcal{H})$ and =$\mathcal{B}=\{B_{k}\}_{k\in K}\subset{\mathcal{B}}(\mathcal{K})$ be two sets of mutually commuting normal operators, and let $A=\{a_{i}\}_{i\in I}\subseteq\mathcal{H}$ and $B=\{b_{i}\}_{i\in I}\subseteq\mathcal{K}$ be two subsets of vectors such that
$\left\langle\left(\prod_{k\in K_{0}}E_{A_{k}}(\Delta_{k})\right)a_{i},a_{j}\right\rangle=\left\langle\left(\prod_{k\in K_{0}}E_{B_{k}}(\Delta_{k})\right)b_{j},b_{i}\right\rangle$
for all finite subsets $K_{0}\subseteq K$, Borel subsets $\{\Delta_{k}\}_{k\in K_{0}}$ of $\mathbb{C}$ and $i,j\in I$.
Then, there exists a conjugate-linear surjective isometry $W:L_{\mathcal{A},A}\to L_{\mathcal{B},B}$ such that
$Wa_{i}=b_{i}\quad\mbox{and}\quad Wp(T_{k_{1}},T_{k_{1}}^{*},\dots,T_{k_{n}},T_{k_{n}}^{*})=p(S_{k_{1}},S_{k_{1}}^{*},\dots,S_{k_{n}},S_{k_{n}}^{*})^{*}W$
for all $i\in I$, $n\in\mathbb{Z}^{+}$, $p\in\mathcal{P}_{2n}(\mathbb{C})$ and $\{k_{l}\}_{1\leq l\leq n}\subseteq K$, where $T_{k_{i}}$ and $S_{k_{i}}$ are the restrictions of $A_{k_{i}}$ and $B_{k_{i}}$ to $L_{\mathcal{A},A}$ and $L_{\mathcal{B},B}$, respectively.
@@
According to Lemma 2.13, for every subset $\{k_{l}\}_{1\leq l\leq n}\subseteq K$, every polynomial $p\in\mathcal{P}_{2n}(\mathbb{C})$, and all $i,j\in I$, we have
(2.2) $\left\langle p(A_{k_{1}},A_{k_{1}}^{*},\dots,A_{k_{n}},A_{k_{n}}^{*})a_{i},a_{j}\right\rangle=\left\langle p(B_{k_{1}},B_{k_{1}}^{*},\dots,B_{k_{n}},B_{k_{n}}^{*})b_{j},b_{i}\right\rangle.$
@@
$\mathcal{H}_{0}={\rm span}\big{\{}p(A_{k_{1}},A_{k_{1}}^{*},\dots,A_{k_{n}},A_{k_{n}}^{*})a_{i}:n\in\mathbb{Z}^{+},p\in\mathcal{P}_{2n}(\mathbb{C}),\{k_{l}\}_{1\leq l\leq n}\subseteq K,i\in I\big{\}}$
@@
$\mathcal{K}_{0}={\rm span}\big{\{}p(B_{k_{1}},B_{k_{1}}^{*},\dots,B_{k_{n}},B_{k_{n}}^{*})b_{i}:n\in\mathbb{Z}^{+},p\in\mathcal{P}_{2n}(\mathbb{C}),\{k_{l}\}_{1\leq l\leq n}\subseteq K,i\in I\big{\}},$
and let $W:\mathcal{H}_{0}\to\mathcal{K}_{0}$ be the conjugate-linear transformation given by
$W\left(p(A_{k_{1}},A_{k_{1}}^{*},\dots,A_{k_{n}},A_{k_{n}}^{*})a_{i}\right)=p(B_{k_{1}},B_{k_{1}}^{*},\dots,B_{k_{n}},B_{k_{n}}^{*})^{*}b_{i}$
for all $n\geq 1$, $p\in\mathcal{P}_{2n}(\mathbb{C})$, $\{k_{l}\}_{1\leq l\leq n}\subseteq K$ and $i\in I$.
@@
We will first show that $W$ is well-defined. Arbitrarily choose $h\in\mathcal{H}_{0}$; then, there exists a finite subset $J$ of $I$, $\{k_{i,l}\}_{1\leq l\leq n_{i}}\subseteq K$, for $i\in J$, and polynomials $p_{i}\in\mathcal{P}_{2n_{i}}(\mathbb{C})$, for $i\in J$, such that
$h=\sum_{i\in J}p_{i}(A_{k_{i,1}},A_{k_{i,1}}^{*},\dots,A_{k_{i,n_{i}}},A_{k_{i,n_{i}}}^{*})a_{i}.$
Then,
$\|h\|^{2}$ $=\left\langle\sum_{i\in J}p_{i}(A_{k_{i,1}},A_{k_{i,1}}^{*},\dots,A_{k_{i,n_{i}}},A_{k_{i,n_{i}}}^{*})a_{i},\sum_{j\in J}p_{j}(A_{k_{j,1}},A_{k_{j,1}}^{*},\dots,A_{k_{j,n_{j}}},A_{k_{j,n_{j}}}^{*})a_{j}\right\rangle$
$=\sum_{i\in J}\sum_{j\in J}\left\langle p_{i}(A_{k_{i,1}},A_{k_{i,1}}^{*},\dots,A_{k_{i,n_{i}}},A_{k_{i,n_{i}}}^{*})a_{i},p_{j}(A_{k_{j,1}},A_{k_{j,1}}^{*},\dots,A_{k_{j,n_{j}}},A_{k_{j,n_{j}}}^{*})a_{j}\right\rangle$
$=\sum_{i\in J}\sum_{j\in J}\left\langle p_{i}(A_{k_{i,1}},A_{k_{i,1}}^{*},\dots,A_{k_{i,n_{i}}},A_{k_{i,n_{i}}}^{*})p_{j}(A_{k_{j,1}},A_{k_{j,1}}^{*},\dots,A_{k_{j,n_{j}}},A_{k_{j,n_{j}}}^{*})^{*}a_{i},a_{j}\right\rangle.$
It follows by (2.2) that
$\|h\|^{2}$ $=\sum_{i\in J}\sum_{j\in J}\left\langle p_{i}(B_{k_{i,1}},B_{k_{i,1}}^{*},\dots,B_{k_{i,n_{i}}},B_{k_{i,n_{i}}}^{*})p_{j}(B_{k_{j,1}},B_{k_{j,1}}^{*},\dots,B_{k_{j,n_{j}}},B_{k_{j,n_{j}}}^{*})^{*}b_{j},b_{i}\right\rangle$
$=\sum_{i\in J}\sum_{j\in J}\left\langle p_{j}(B_{k_{j,1}},B_{k_{j,1}}^{*},\dots,B_{k_{j,n_{j}}},B_{k_{j,n_{j}}}^{*})^{*}b_{j},p_{i}(B_{k_{i,1}},B_{k_{i,1}}^{*},\dots,B_{k_{i,n_{i}}},B_{k_{i,n_{i}}}^{*})^{*}b_{i}\right\rangle$
$=\left\langle\sum_{j\in J}p_{j}(B_{k_{j,1}},B_{k_{j,1}}^{*},\dots,B_{k_{j,n_{j}}},B_{k_{j,n_{j}}}^{*})^{*}b_{j},\sum_{i\in J}p_{i}(B_{k_{i,1}},B_{k_{i,1}}^{*},\dots,B_{k_{i,n_{i}}},B_{k_{i,n_{i}}}^{*})^{*}b_{i}\right\rangle$
$=\|Wh\|^{2}.$
@@
Hence, if $h,k\in\mathcal{H}_{0}$ are such that $\|h-k\|=0$, we get $\|Wh-Wk\|=\|W(h-k)\|=\|h-k\|=0$. Therefore, $W$ is a well-defined conjugate-linear isometry. Moreover, by its construction, $W$ is surjective. By density of $\mathcal{H}_{0}$ and $\mathcal{K}_{0}$ in $L_{\mathcal{A},A}$ and $L_{\mathcal{B},B}$, respectively, it follows that $W$ can be extended to a conjugate-linear surjective isometry from $L_{\mathcal{A},A}$ to $L_{\mathcal{B},B}$, which we also denote by $W$. Furthermore, it can be seen that $Wa_{i}=b_{i}$ for every $i\in I$.
Now, let $n\in\mathbb{Z}^{+}$, $p\in\mathcal{P}_{2n}(\mathbb{C})$ and $\{k_{l}\}_{1\leq l\leq n}\subseteq K$, and let us check that
$T:=Wp(T_{k_{1}},T_{k_{1}}^{*},\dots,T_{k_{n}},T_{k_{n}}^{*})=S:=p(S_{k_{1}},S_{k_{1}}^{*},\dots,S_{k_{n}},S_{k_{n}}^{*})^{*}W.$
Let $m\in\mathbb{Z}^{+}$, $q\in\mathcal{P}_{2m}(\mathbb{C})$, $\{t_{r}\}_{1\leq r\leq m}\subseteq K$ and $i\in I$, and put $h=q(A_{t_{1}},A_{t_{1}}^{*},\dots,A_{t_{m}},A_{t_{m}}^{*})a_{i}$. Then,
$Th$ $=Wp(T_{k_{1}},T_{k_{1}}^{*},\dots,T_{k_{n}},T_{k_{n}}^{*})h$
$=Wp(A_{k_{1}},A_{k_{1}}^{*},\dots,A_{k_{n}},A_{k_{n}}^{*})h$
$=Wp(A_{k_{1}},A_{k_{1}}^{*},\dots,A_{k_{n}},A_{k_{n}}^{*})q(A_{t_{1}},A_{t_{1}}^{*},\dots,A_{t_{m}},A_{t_{m}}^{*})a_{i}$
$=q(B_{t_{1}},B_{t_{1}}^{*},\dots,B_{t_{m}},B_{t_{m}}^{*})^{*}p(B_{k_{1}},B_{k_{1}}^{*},\dots,B_{k_{n}},B_{k_{n}}^{*})^{*}b_{i}$
$=p(B_{k_{1}},B_{k_{1}}^{*},\dots,B_{k_{n}},B_{k_{n}}^{*})^{*}q(B_{t_{1}},B_{t_{1}}^{*},\dots,B_{t_{m}},B_{t_{m}}^{*})^{*}b_{i}$
$=p(S_{k_{1}},S_{k_{1}}^{*},\dots,S_{k_{n}},S_{k_{n}}^{*})^{*}Wh=Sh.$